OVERTONE ABSORPTIONS OF METHANE: ANOMALOUS TRANSITION FROM NORMAL MODE TO LOCAL MODE

\underline{C. MANZANARES}, D. L. CEDENO and V. M. BLUNT, Department of Chemistry, Baylor University, Waco, TX 76798.